# Optimal RIS Placement in a Multi-User MISO System with User Randomness

### v1

@@ IV-B The Outer Optimization Problem
Since closed-form solutions for $\mathbf{w}$ and $\varphi$ as functions of the channel, valid for all possible instantiations of the user distribution is clearly intractable, we approximate the expectation in $\mathbf{P1.1}$ over multiple realizations of the user distribution.
For each instantiation of a set of users, a discrete set of candidate locations for RIS (of cardinality $T$) is formed using Algorithm 1 assuming that Circle $C$ is the entire cell. The main idea behind the algorithm stems from the necessity for the RIS to have a clear Line of Sight from the BS to be able to forward the message to the user. In addition, the RIS is placed at least $d_{FF}=\frac{2D^{2}}{\lambda_{c}}$ far from the BS (Fraunhofer distance of the BS), so that the RIS does not interfere in the near field of the BS [23]. Here, $\lambda_{c}$ denotes the wavelength corresponding to the frequency of operation and $D$ denotes the overall aperture size of the BS.
@@
3:  Generate random $\mathbf{q}_{i}$ in $(i\text{ mod }4)^{th}$ quadrant of $C$
4:  if No obstacle lies on the direct link from $\mathbf{o_{b}}$ to $\mathbf{q}_{i}$ and $\mathbf{q_{i}}$ is at least $d_{FF}$ away then
@@
A method to construct a set of feasible ‘min-max’ solutions $\mathcal{S}$ is proposed in Algorithm 2 based on Algorithm 1.
@@
1:Create multiple instantiations of user distribution and use Algorithm 1 to get corresponding candidate sets in $C$
2:For each candidate in each of these candidate sets corresponding to each instantiation, perform joint beamforming
3:For each instantiation, a feasible solution is obtained using $\mathbf{P1.1}$, forming a solution set $\mathcal{S}$
IV-C Computing the Final Solution
It is important to observe that the solution set is not closed under linear combinations, i.e. $s_{1}\in\mathcal{S}$ and $s_{2}\in\mathcal{S}\nRightarrow(a.s_{1}+b.s_{2})\in\mathcal{S},\forall a,b\in\mathbb{R}$. This rules out the usual tendency to compute the mean of all optimal locations to find the ‘final solution’ (mean of solutions not necessarily a solution itself).
Algorithm 3 is a recursive method proposed to compute the final solution. The algorithm was developed by observing the spatial distribution of the solutions for a wide variety of scenarios and user distributions. ‘Clusters’ are formed among the solutions in $\mathcal{S}$ by creating quantized versions of them with a suitable step size, $d_{start}$. Then the cluster with the highest number of entries is declared to be the ‘optimal cluster’. Then Algorithm 2 is repeated for a new candidate set with candidates only within the optimal cluster. This process of looking deeper into the optimal cluster is repeated up to the required precision $d_{p}$. Hence, the globally optimal region to deploy the RIS is obtained as the square $\mathbf{\hat{x}_{r}}(1)-d_{p}/2,\mathbf{\hat{x}_{r}}(1)+d_{p}/2)\times[\mathbf{\hat{x}_{r}}(2)-d_{p}/2,\mathbf{\hat{x}_{r}}(2)+d_{p}/2)$.
@@ V Simulation Results
(a) Scenario 1: Solutions - RL 1.
(b) Scenario 1: Frequency - RL 1.
(c) Scenario 1: Frequency - RL 2.
(d) Scenario 1: Final Solution.
(e) Scenario 2: Final Solution.
(f) Scenario 3: Final Solution.
(g) Scenario 4: Final Solution.
(h) Coverage Analysis.
Figure 2: Optimal RIS placement in different scenarios.
@@ V-A Simulation Setup
where $d_{i}$ represents the Euclidean distance traversed by each link in meters (m) and $f_{c}$ is the carrier frequency in GHz. $f_{c}$ is set as $2.4$ GHz. For example, we get a loss of $\beta_{br}=\beta_{ru_{k}}=55.6$ dB and $\beta_{bu_{k}}=69.3$ dB for $d_{br}=d_{ru_{k}}=d_{bu_{k}}=10$ m. Then the Rician factors are set to $T_{1}=T_{2}=10$ dB. The transmit power is set to $P_{max}=0$ dB unless mentioned otherwise. In addition,
$\sigma^{2}=-174+\log_{10}(B)+NF\quad\text{ dBm},$ (16)
where bandwidth is set to $B=10$ MHz (since we consider single-carrier communication) and receiver noise figure $NF=5$ dB [24]. Also, in this work, we assume that when a link hits an obstacle, the link power is reduced to a negligible level due to the severe penetration loss.
@@ V-B Results and Discussion
Four different scenarios are considered for simulation as shown in Fig. 2. The first 3 scenarios assume various positions, sizes, and orientations for obstacles that hinder access to a set of users following a homogeneous PPP with a fixed density $\lambda_{u}$ throughout the circular cell. And finally, a more practical scenario with 4 user hotspots is considered.
@@
Fig. 2(a)-(d) present a top-level view of the complete procedure described in Section IV for Scenario 1. Fig. 2(a) shows a scatterplot of the solution set in recursion level (RL) 1, in which the black objects are the obstacles, the green circle represents the user cluster (the whole cell), and the blue dots represent the solutions. Fig. 2(b) represents the frequency heatmap of a quantized version of this solution set with $d_{start}=2$ m. To achieve $d_{p}=1\text{ m}$, another recursive pass of the algorithm is run as shown in Fig. 2(c). The cluster with the highest RL 2 frequency is declared as the optimal region for RIS deployment (represented by a $d_{p}\times d_{p}$ square) in Fig. 2(d). As shown in Fig. 2(h), deploying RIS at this location extends coverage from 65.76% to 99.68% of the cell. The average WSR of deploying RIS at the center of this cluster is captured in Table I. Note that the expected number of users corresponding to this scenario is $E[K]=\lambda_{u}.\text{Area}=11.3097$.
TABLE I: Obtained results
Scenarios $\mathbf{\lambda_{u}}$ Final Solution Average WSR Coverage
(bps/Hz)
1 0.009 (105 m,31 m) 5.5631 99.68%
2 0.007 (105 m,42 m) 5.8089 97.91%
3 0.009 (113 m,31 m) 5.8040 97.21%
4 $\begin{cases}0.4,&\text{Circle }C_{1}\\ 0.04,&\text{Circle }C_{2}\\ 0.05,&\text{Circle }C_{3}\\ 0.035,&\text{Circle }C_{4}\\ 0,&\text{elsewhere}\end{cases}$ (111 m,41 m) 3.5982 99.69%
@@
A major obstacle group is considered at the center of the cell in Scenario 2, as shown in Fig. 2(e). A lower user density (on average, three fewer users) is assumed. The optimal location and average WSR is given in Table I. As shown in Fig. 2(h), this extends coverage from 66.27% to 97.91% of the cell. Scenario 3 as shown in Fig. 2(f) was designed to be difficult to extend coverage to the entire cell with a single RIS. The algorithm still derives a solution that covers regions shadowed by the major clusters and covers a portion of the region shadowed by the single wall obstacle (i.e., extends coverage to 97.21% of the cell), with a similar WSR performance.
@@
Finally, a more practical Scenario 4 is considered as depicted in Fig. 2(g) with location dependent $\lambda_{u}$ as given in Table I. Here, Circles $C_{1},C_{2},C_{3},C_{4}$ are assumed as shown in Fig. 2(h). The same obstacle configuration as in Scenario 3 is assumed. The final solution shifts from (109 m, 45 m) to (114 m, 42 m) to accommodate the significantly higher user density in the region shadowed by the wall obstacle, highlighting the capability of the proposed method to converge to a location that not only maximizes coverage by covering most of the shadowed regions, but also effectively adapts to spatial variations in user density.
Figure 3: WSR-transmit power characteristic for Scenario 2.
The average WSR vs transmit power characteristics for Scenario 2 is shown in Fig. 3. The average WSR of deploying RIS at the optimal location significantly exceeds the ‘No RIS’ simulation. This is the increase achieved solely because of the extension of coverage by the RIS to almost the full cell. Furthermore, when the RIS is randomly positioned within the cell, and for each of these random locations - averaged over multiple realizations of the user distribution, the resulting average performance remains close to that of the ‘No RIS’ scenario. This observation underscores the necessity for the optimal placement of RIS. The range of WSR performance across scenarios aligns with similar works [11, 12], reaffirming the optimality of the solution obtained in our work. It is also important to note that this is achieved despite having lower $M$ ($16<32$), practical number of users (average $K\approx 11\gg 2$), different obstacle configurations and the randomness of users.
@@ VI Conclusion
A novel method was proposed to tackle typical practical requirements of an RIS-assisted system, such as exploiting the user distribution to maximize performance, exploiting the obstacle configuration to maximize the coverage of the system, and handling user randomness. This is the first work, to the best of our knowledge, to handle the ideal hierarchical optimization directly - maximizing coverage, while also ensuring good performance. A set of min-max solutions for the RIS placement was proposed to be obtained by searching over multiple discrete sets formed based on numerous realizations from the user distribution. Finally, a recursive strategy was used to find the final region to deploy the RIS. The superiority of the proposed algorithm was validated using simulations.

### v2

@@ IV-B The Outer Optimization Problem
Since closed-form solutions for $\mathbf{w}$ and $\varphi$ as functions of the channel in $\mathbf{P1.2}$ is clearly intractable, while solving $\mathbf{P1.1}$, we are instead required to approximate the expectation of SINR over multiple realizations of the user distribution.
A discrete set of candidate locations for RIS (of cardinality $T$) can be formed using Algorithm 1 assuming that Circle $C$ to be the entire cell. The main idea behind the algorithm stems from the necessity for the RIS to have a clear Line of Sight (LoS) from the BS to be able to forward the message to all users. In addition, the RIS is placed at least $d_{FF}=\frac{2D^{2}}{\lambda_{c}}$ far from the BS (Fraunhofer distance of the BS) [25], so that the RIS does not interfere in the near field of the BS. Here, $\lambda_{c}$ denotes the wavelength corresponding to the frequency of operation and $D$ denotes the overall aperture size of the BS.
@@
3:  Pick a random $\mathbf{q}_{i}$ in $C$
4:  if $\mathbf{q_{i}}$ is at least $d_{FF}$ away from $\mathbf{o_{b}}$ and Clear LoS from $\mathbf{o_{b}}$ to $\mathbf{q}_{i}$ then
@@
Furthermore, a set of feasible ‘max-min’ solutions (set $\mathcal{S}$) can be constructed using these candidate sets as outlined in Algorithm 2. For a large number of instantiations of sets of users from the homogeneous PPP distribution, we first form corresponding candidate sets using Algorithm 1. Then, joint beamforming is performed assuming that the RIS is placed at each candidate location in each of the candidate sets (one after another). Finally, from each candidate set, that RIS that maximizes the performance of the worst performing user in that set of $K$ users, is added as a solution to $\mathcal{S}$.
@@
1:Create multiple instantiations of users from its distribution and use Algorithm 1 to get corresponding candidate sets
2:For each candidate, in each candidate set, corresponding to an instantiation of users, perform joint beamforming
3:For each instantiation of a set of users, a feasible solution is obtained using $\mathbf{P1.1}$, forming a solution set $\mathcal{S}$
@@
IV-C Recursive Clustering
It is important to observe that this solution set is not closed under linear combinations, i.e. $s_{1}\in\mathcal{S}$ and $s_{2}\in\mathcal{S}\nRightarrow(a.s_{1}+b.s_{2})\in\mathcal{S},\forall a,b\in\mathbb{R}$. This rules out the usual tendency to compute the mean of all optimal locations in $\mathcal{S}$ to find the ‘final solution’ (mean of solutions not necessarily a solution itself).
TABLE I: Time complexity comparison with benchmarks
Algorithm Complexity of $\mathbf{W}$ update Complexity of $\boldsymbol{\theta}$ update Overall Complexity Order
Proposed $\mathcal{O}(KM^{2})$ $\mathcal{O}(KN^{2}\min{\{M,K\}})$ $\mathcal{O}(I_{\text{iter}}(KM^{2}+KN^{2}\min{\{M,K\}}))$
[12] $\mathcal{O}(K(2M^{3}+NM^{2}))$ $\mathcal{O}(IG(NM^{2}+MN+M))$ $\mathcal{O}(I_{\text{iter}}(M^{3}+MN^{2}+K(2M^{3}+NM^{2})+IG(NM^{2}+MN+M)+KMN)$
[13] $O(K^{2}V)$ $O(K^{2}JW)$ $\mathcal{O}(I_{\text{iter}}(K^{2}V+K^{2}JW))$
Therefore, we propose Algorithm 3 which is a recursive method to compute the final solution. ‘Clusters’ are formed among the solutions in $\mathcal{S}$ by creating quantized versions of them with a suitable step size, $d_{start}$. Then the cluster with the highest number of entries is declared to be the ‘optimal cluster’. Then Algorithm 2 is repeated for a new candidate set with candidates only within the optimal cluster. This process of looking deeper into the optimal cluster is repeated up to the required precision $d_{p}$. Hence, the globally optimal region to deploy the RIS is obtained as the square $\mathbf{\hat{x}_{r}}(1)-d_{p}/2,\mathbf{\hat{x}_{r}}(1)+d_{p}/2)\times[\mathbf{\hat{x}_{r}}(2)-d_{p}/2,\mathbf{\hat{x}_{r}}(2)+d_{p}/2)$.
Remark 1
The core part of the proposed algorithm only requires the obstacle configuration, channel parameters, and the user distribution. Owing to the separable structure of the problem formulation in $\mathbf{P1.1}$ and $\mathbf{P1.2}$, the algorithm is capable of using any proposed or future joint beamforming technique. And there are many existing joint beamforming algorithms that work under imperfect CSI conditions including [20]. In our work, we have used our adaptation of a well-accepted beamforming algorithm [18] that works with perfect CSI. That clearly does not limit the proposed algorithm in any way to work only under perfect CSI conditions. In other words, the proposed approach remains effective for any CSI acquisition and RIS training technique with the corresponding channel estimation errors, which jointly affect the SINR expression in Eq. (8), prior to its use in our algorithm.
IV-D Time Complexity and Efficient Implementation
The overall computational efficiency of the proposed algorithm is strongly dependent on the efficiency of the inner optimization. The time complexity of updating $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ using Eq. (12) and (13) can be easily shown to be $\mathcal{O}(KMN)$.
On the other hand, optimization of transmit beamforming from Eq. (14) involves matrix inversion, making this step $\mathcal{O}(KMN+I_{\lambda}M^{3})$ where $I_{\lambda}$ is the number of iterations for the bisection search for $\lambda$. This step could potentially become a bottleneck for the efficiency of the proposed solution. Hence, we adopt a linear update [18]
$\mathbf{w}_{k}=\frac{1}{L-\lambda}(L\hat{\mathbf{w}}_{k}-\mathbf{g}_{k}),$ (20)
[…]
where $\hat{\mathbf{w}}_{k}=\mathbf{w}_{k}+\epsilon(\mathbf{w}_{k}-\mathbf{w}^{\text{prev}}_{k})$ for some $\epsilon>0$. Here, $\mathbf{g}_{k}=-\frac{\partial f_{2}}{\partial\mathbf{w}_{k}}$ and $L$ is set to the Lipschitz constant of $\mathbf{g}_{k}$. This linear update results in the same WMMSE solution in Eq. (14), but reduces time complexity to a reasonable $\mathcal{O}(KM^{2})$.
Similarly, the phase optimization is dominated by the calculation of $\boldsymbol{U}$ and the gradient projection algorithm. The resulting time complexity of the direct implementation is $\mathcal{O}(K^{2}N^{2}+I_{\rho}N^{2})=\mathcal{O}(K^{2}N^{2})$ where $I_{\rho}$ is the number of iterations spent in calculating the dynamic step size $\rho$. Instead of implementing $\boldsymbol{U}$ directly like [18] or [12], one can also take the summation $\sum_{i=1}^{K}\mathbf{w}_{i}\mathbf{w}^{H}_{i}$ outside and compute it just once to get a time complexity of $\mathcal{O}(KMN^{2})$. Since $K$ is a Poisson random variable in our implementation, we adopt a dynamic strategy to use the direct implementation of $\boldsymbol{U}$ if it is the usual case of $M\geq K$ and the second if $M<K$. This results in the best complexity of $\mathcal{O}(KN^{2}\min\{K,M\})$.
Hence, the overall complexity of the proposed solution is $\mathcal{O}(I_{RL}\times I_{AO}\times(KMN+KM^{2}+KN^{2}\min\{K,M\}))$ where $I_{RL}$ is the number of recursive levels traced in the outer optimization and $I_{AO}$ is the number of inner optimization iterations. It is also important to note that this is a one time cost spent before the RIS assisted system is actually implemented.
A comparison of the computational efficiency of our method with similar benchmarks is provided in Table I. Let $I_{\text{iter}}$ uniformly denote the total number of iterations we execute the considered algorithm. For our algorithm, $I_{\text{iter}}=I_{RL}\times I_{AO}$. From Table I, [12] uses the matrix inversion implementation for transmit beamforming. Also, it assumes that the number of users and their actual locations are precisely known before placement which is not practical. Our implementation is more efficient and clearly more practical. On the other hand, the time complexity of the analog beamforming in [13] is comparable to that of the beamformer update in our proposed method. The time complexity of $\mathcal{O}(K^{2}JW)$ for phase optimization is very similar to the complexity of direct implementation, $\mathcal{O}(K^{2}N^{2})$. Ours is a more generalized version of the same order.
TABLE II: Optimal RIS Placement in different scenarios
Scenarios $\mathbf{\lambda_{u}}$ Average WSR Coverage (%)
(bps/Hz) With RIS Without RIS
1 0.009 6.7915 99.68% 65.76%
2 0.007 7.6556 97.91% 66.27%
3 0.009 6.7522 97.21% 75.96%
4 $\begin{cases}0.4,&\text{Circle }C_{1}\\ 0.04,&\text{Circle }C_{2}\\ 0.05,&\text{Circle }C_{3}\\ 0.035,&\text{Circle }C_{4}\\ 0,&\text{elsewhere}\end{cases}$ 6.6982 99.69% 70.17%
(a) Scenario 1: Final Solution.
(b) Scenario 2: Final Solution.
(c) Scenario 3: Final Solution.
(d) Scenario 4: Final Solution.
Figure 2: Optimal RIS placement in different scenarios.
(a) Set of min-max solutions obtained in Recursion Level 1.
(b) Frequency map of the quantized solution set in Recursion Level 1.
(c) Frequency map of the quantized solution set in Recursion Level 2.
Figure 3: Recursive Clustering for Optimal RIS placement in Scenario 1.
Remark 2
Since $K$ in our system is random, all analytical results are derived in general ($K$ may be greater than, equal to, or less than the number of transmit antennas $M$). However, it is important to note that when the system is overloaded ($M<K$), the derived WMMSE result is no longer optimal unless it is combined with an appropriate scheduling algorithm, which will select $S\leq M$ users that jointly employs beamforming and hence do not interfere with each other. For scheduling purposes, these $S$ users are treated as a single “virtual” user occupying one time slot. The remaining KSK-S users may interfere with each other and with this virtual user. So, their transmission is handled via orthogonal multiple access (to mitigate the interference). Although for simplicity, we assume that $M\geq K$ in the simulations presented, we emphasize that the proposed algorithm is general and can be effectively integrated with any scheduling algorithm in literature. A detailed investigation is beyond the scope of this paper and is omitted for brevity and space limits.
@@ V-A Simulation Setup
where $d_{i}$ represents the Euclidean distance traversed by each link in meters (m) and $f_{c}$ is the carrier frequency in GHz. $f_{c}$ is set as $2.4$ GHz. For example, we get a loss of $\beta_{br}=\beta_{ru_{k}}=55.6$ dB and $\beta_{bu_{k}}=69.3$ dB for $d_{br}=d_{ru_{k}}=d_{bu_{k}}=10$ m. Then the Rician factors are set to $T_{1}=T_{2}=10$ dB. The transmit power is set to $P_{max}=0$ dB and $\sigma^{2}=-81$ dBm [4] unless mentioned otherwise. We also assume that when a link hits an obstacle, the link power is reduced to a negligible level due to the severe penetration loss.
@@ V-B Optimal RIS Placement
Four different scenarios are considered for simulation as shown in Table II and Fig. 2. The first three scenarios assume various positions, sizes, and orientations for obstacles that hinder access to a set of users following a homogeneous PPP with a fixed density $\lambda_{u}$ throughout the circular cell. Finally, a more practical scenario with four user hotspots is considered.
@@
Fig. 3 (a)-(c) present a top-level view of the complete procedure described in Section IV for Scenario 1. Fig. 3 (a) shows a scatterplot of the solution set in recursion level (RL) 1, in which the black objects are the obstacles, the green circle represents the user cluster (the whole cell in this case), and the blue dots represent the solutions. Fig. 3 (b) represents the frequency heatmap of a quantized version of this solution set with $d_{start}=2$ m. To achieve $d_{p}=1$ m, another recursive pass of the algorithm is run as shown in Fig. 3 (c). The cluster with the highest RL 2 frequency is declared as the optimal region for RIS deployment (represented by a $d_{p}\times d_{p}$ square) in Fig. 2 (a). As per Table II, deploying RIS at this location extends coverage from 65.76% to 99.68% of the cell under consideration. The average WSR performance of deploying RIS at the center of this cluster is also captured in Table II. Note that the expected number of users corresponding to this scenario is a practical $E[K]=\lambda_{u}.\text{Area}=11.3097$.
(a) $P_{max}$ characteristic.
(b) Convergence analysis.
Figure 4: Conventional Performance analysis of proposed algorithm in Scenario 2.
(a) RIS positional analysis.
(b) Obstacle analysis.
Figure 5: Further analysis of proposed algorithm for Scenario 2.
@@
A major obstacle group is considered at the center of the cell in Scenario 2, as shown in Fig. 2 (b). A lower user density (on average, three fewer users) is assumed. The optimal location and average WSR are given in Table II. As shown in Table II, this extends coverage from 66.27% to 97.91% of the cell. Scenario 3, as shown in Fig. 2 (c), was designed to be difficult to extend coverage to the entire cell with a single RIS. The algorithm still derives a solution that covers regions shadowed by the major clusters and covers a portion of the region shadowed by the single wall obstacle (i.e., extends coverage to 97.21% of the cell), with a similar WSR performance.
@@
Finally, a practical Scenario 4 with four user hotspots at $C_{1},C_{2},C_{3},C_{4}$ is considered as depicted in Fig. 2 (d). The scenario also assumes a spatially dependent $\lambda_{u}$, as given in Table II. Furthermore, although this scenario assumes the same obstacle configuration as in Scenario 3, the final solution shifts from (109 m, 45 m) to (114 m, 42 m) to accommodate the increase in user density in the region shadowed by the wall obstacle, highlighting the capability of the proposed method to converge to a location that not only maximizes coverage by covering most of the shadowed regions, but also effectively adapts to spatial variations in user density, resulting in improved WSR performance.
V-C Performance Analysis
The average WSR vs. transmit power characteristics for Scenario 2 is shown in Fig. 4 (a). The performance of the proposed algorithm is much greater than the ‘without RIS’ and the ‘random phase’ cases, reaffirming the importance of an RIS in the system. In addition, the average WSR for deploying RIS in the optimal location clearly exceeds the ‘Random Placement’ simulation. This gain proves that optimal placement of RIS, not only results in an extension of coverage to about 98% (or increase in fairness as per our objective $\mathbf{P1.1}$), but also results in an increase in the performance of the system. Furthermore, the WSR performance of the proposed scheme exceeds the performance of similar works [12, 13]. For instance, when $P_{max}=40$ dBm, a gain of 3.4 and 2.7 bps/Hz is achieved over a multi-RIS assisted system that employs location optimization and hybrid beamforming for an indoor setup [13] and a co-located multi-RIS assisted MISO system that employs joint location and beamforming optimization [12], respectively. It is also important to note that, owing to the optimality of the proposed scheme, this gain is achieved despite having a lower $M=16<32$, just a single RIS, practical number of users (average $K\approx 11\gg 2$), different obstacle configurations, and the randomness of users.
Fig. 4 (b) provides the convergence analysis of the inner optimization. For instance, the average WSR performance for $5$ AO iterations is obtained by averaging over multiple user instantiations, where the performance of each instantiation is evaluated by performing joint beamforming for 5 iterations.
Then, Fig. 5 (a) provides a detailed performance analysis of shifting the RIS across the vertical diameter of the cell in Scenario 2. As depicted, the converged solution (105 m, 42 m) maximizes both coverage and average WSR as desired.
Finally, Fig. 5 (b) generalizes Scenario 2 by varying the number of obstacles in the central cluster. As anticipated, performance degrades with the increase in number of obstacles. However, the rate of fall in performance of the proposed algorithm is much lesser than that of ‘random placement’ due to the optimality of the proposed algorithm. Also, notably, when the number of obstacles in the central cluster is 17 or more (placed across the cell’s vertical diameter), there is no way an RIS can help improve performance, forcing the proposed algorithm to perform like random placement.
@@ VI Conclusion
A novel method was proposed to tackle typical practical requirements of an RIS-assisted system, such as exploiting the user distribution to maximize performance, exploiting the obstacle configuration to maximize the coverage of the system, and handling user randomness. This is the first work, to the best of our knowledge, to handle the ideal hierarchical optimization directly - maximizing coverage, while also ensuring good performance. It is also notable that the method is valid for both perfect and imperfect CSI conditions depending on the adopted beamforming algorithm. A set of min-max solutions for the RIS placement was proposed to be obtained by searching over multiple discrete sets formed based on numerous realizations from the user distribution. Finally, a recursive strategy was used to find the final region to deploy the RIS. The superiority of the proposed algorithm was validated using simulations.
Notably, the proposed algorithm, both inner and outer optimizations, can be readily modified for a Multi-RIS assisted system. We look forward to investigating the same in our future work.